is 1degree equal to 1radian
 

Grade:10

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
radians and degrees are both units of measurement of the unit circle.
1 degree is not equal to 1 radian
radians = (degrees x\pi) / 180
degrees = (radians x 180) /\pi
so 1 radian=(\pi)/180
Thanks and Regards
Shaik Aasif
askIITians faculty